OpenStudy (boxman61):

--Answer Check-- A particular kind of cork has a density of 0.30 g/mL , so it floats. Could you lift a cork ball one meter in diameter?

OpenStudy (boxman61):

\[V=\frac{ 4 }{ 3 }\pi r^3\] 1m=39.4in 1in=2.54cm^3 D=39.4in R=19.7in so Volume= 32,023in^3 32,023in^3 * 2.54cm^3 = 81,341cm^3 Volume 81,341cm^3 = 81,341mL 81,341mL * 0.30g/mL = 24,402.3g or 53.7lbs So yes hopefully you could lift the cork ball

OpenStudy (welshfella):

Not sure why you converted to inches and lbs Volume of a cork ball 1 m in diameter = (4/3)*pi* 50^3 cms^3 its mass = (4/3)*pi* 50^3 * 0.3 = 157079.6 gms = 157.08 kg.

OpenStudy (welshfella):

Most people couldn't lift that unless they are a weightlifter

OpenStudy (boxman61):

Quick question, where did the Radius of 50 come from?